Can someone help? This problem has me really torn, I can't seem to solve it. Any help would be great. Perpetual Inventory Using FIFO : Beginning inventory, purchases, and sales for Item ER27 are as follows: Nov. 1 Inventory 40Use the Dynamic Exhibit to answer the following questions. 1. When the number of units sold on November 21 is 36, the cost of
Answer #1

1) Cost of merchandise sold on November 21 = (10*5+26*7) = 232

2) Ending inventory = 44*7 = $308

3) cost of merchandise sold on November 21 = 8*5 = 40

4) Ending inventory = 30*7 = 210